November 1996  ICTCM IX (Reno)  11

Step 1 – Free-Fall Solution

> SOL1 := dsolve( subs(k=k1,SYS) union IC, { x(t), v(t) } ):
> V1 := subs( SOL1, v(t) );

V1 := − \frac{m g}{k1} + \frac{m g e^{(− \frac{t k1}{m} )}}{k1}

> X1 := subs( SOL1, x(t) );

X1 := \frac{−m g t k1 + m^{2} g + x0 k1^{2} − m^{2} g e^{(− \frac{t k1}{m} )}}{k1^{2}}